# Online Learning of Interaction Dynamics with Dual Model Predictive Control for Multi-Agent Systems Using Gaussian Processes

### v1

@@ V Results
We compare our proposed learning GP-based dual MPC (GP-MPC), with a stochastic baseline MPC that uses a constant velocity prediction (CV-MPC) for the Follower, and evaluate their performance on the lane merging use case detailed in Section IV. We test the GP-MPC with and without pre-training. Here, we refer to pre-training as the availability of past measurement data at the start of the experiment.
TABLE II: Results of the Monte Carlo simulations.
@@
V-A Numerical Case Study
[…]
Table II summarizes the results of the Monte Carlo trials. In this study, we consider a merge successful if the Ego merges between the Follower and Leader. The prediction error ${\|e\|\left(k\right):=\frac{1}{N}\|\{v_{i\mid k}-v_{k+i}\}_{i=1}^{N}\|_{1}}$ is the normalized norm of the difference between the predicted trajectory at time $k$ and the realized trajectory from time $k$ to $k+N$. The average prediction error over all time steps and all trials $\overline{\|e\|}$ quantifies the prediction quality of the CV-MPC and GP-MPC.
@@ V-B Discussion
The GP-MPC jointly optimizes the predictions and the associated uncertainty through the dual control effect, as seen in Fig. 2, which shows a successful merge by the GP-MPC without pre-training. Here, the two transparent rectangles indicate the $2\sigma$-bounds on the final predicted position of the Follower. The approximate covariance of the GP is exploited to account for uncertain interactions providing inherent caution through dual MPC. As the Ego is agnostic to the Follower’s response to the merging action it takes extra caution when merging by further tightening the constraints.
Figure 2: Time lapse of a lane merge by the GP-MPC.
Both the CV-MPC and the GP-MPC without pre-training do not capture the high frequent dynamics during the merge, as seen in Fig. 3. During the merge, approximately after $t=10$ [s], the GP-MPC without pre-training has an increased prediction error as it has not observed these high frequent dynamics. Still, the GP-MPC has superior predictions quality prior to merging and is able to identify a safe gap to merge in between. The GP-MPC with pre-training exploits past observations from a similar scenario, leveraging Bayesian inference and the dual control effect to exercise less caution and shows increased prediction quality. The online learning GP-MPC sees a $25\%$ to $29\%$ increase in successful merges compared to CV-MPC, with and without pre-training, respectively. While the complexity of evaluating the sparse GP is linear in the number of data points [15], which is reflected by the average computation time $\overline{T}_{c}$ (Table II), online learning can reduce the required amount of training data. Simulation results show that our GP-MPC can successfully predict the Follower’s motion from various initial conditions by only relying on online training data, demonstrating its generalizability beyond a priori available training data.
@@ VI Conclusions
We presented a Gaussian process-based MPC that uses the dual control effect to learn and predict the dynamics of other agents while accounting for their joint uncertainty. We demonstrated our GP-MPC in a numerical case study on autonomous lane merging, by leveraging dual MPC to jointly optimize the motion of the ego vehicle and the predictions of a target vehicle through an interactive GP model. Through online learning, the GP-MPC can adapt to uncertain and unseen behavior of other agents, showing strong potential for scenarios that extend beyond a priori available training data sets. By explicitly considering the effect of control inputs on the covariance of predictions of other agents, our GP-MPC can leverage this uncertainty in the tightening of constraints.
The dual control effect is an important notion that can be leveraged by the MPC. While this paper focuses on a proof of concept of the proposed method, the next research steps for learning-based GP-MPC include: (i) incorporating the probing effect of dual control, i.e., active learning, (ii) predicting the residual dynamics of multiple agents, (iii) certification and validation of the prediction model, and (iv) providing performance and recursive feasibility guarantees of the MPC.
Figure 3: Average prediction error of Follower’s velocity in experiments in which all planners merge behind (top) and merge in between (bottom).

### v2

@@ V Results
We compare our proposed GP learning-based dual MPC (GP-MPC) with a stochastic baseline MPC that uses a constant velocity prediction (CV-MPC) for the Follower, and evaluate their performance on the lane merging use case detailed in Sec. IV. We test the GP-MPC with and without pre-training. Here, we refer to pre-training as the availability of past measurement data at the start of the experiment.
V-A Numerical Case Study
[…]
TABLE I: Results of the Monte Carlo simulations.
@@ V-A Numerical Case Study
Table I summarizes the results of the Monte Carlo trials. In this study, we consider a merge successful if the Ego merges between the Follower and Leader. The prediction error ${\|e\left(k\right)\|:=\frac{1}{N}\|\{v_{i\mid k}-v_{k+i}\}_{i=1}^{N}\|_{1}}$ is the normalized norm of the difference between the predicted trajectory at time $k$ and the realized trajectory from time $k$ to $k+N$. The average prediction error over all time steps and all trials $\overline{\|e\|}$ quantifies the prediction quality of the CV-MPC and GP-MPC. The average solve time $\bar{T}_{c}$ and the percentage of iterations within sampling time $T_{c}<T_{s}$ show that the algorithm can be run (near-)realtime with these amounts of training data.
@@ V-B Discussion
The GP-MPC jointly optimizes the predictions and the associated uncertainty through the dual control effect, as seen in Fig. 1, which shows a successful merge by the GP-MPC without pre-training. The approximate covariance of the GP is exploited to account for uncertain interactions providing inherent caution through dual MPC. As the Ego is agnostic to the Follower’s response to the merging action, it takes extra caution when merging by further tightening the constraints.
Both the CV-MPC and the GP-MPC without pre-training do not capture the high frequent dynamics during the merge, as seen in Fig. 2. During the merge, approximately after $t=10$ [s], the GP-MPC without pre-training has an increased prediction error as it has not observed these high frequent dynamics. Still, the GP-MPC has superior prediction quality prior to merging and is able to identify a safe gap to merge in between. The GP-MPC with pre-training exploits past observations from a similar scenario, leveraging Bayesian inference and the dual control effect to exercise less caution and shows increased prediction quality. When compared to CV-MPC, the online learning GP-MPC sees a $25\%$ to $29\%$ increase in successful merges over all experiments without and with pre-training, respectively, cf. Table I. While GPs are attractive for Bayesian inference, the computational complexity increases with the number of training data. This stresses the importance of online learning in dynamic environments, such that the MPC does not have to rely on large sets of training data, but rather adapt online. Simulation results show that our GP-MPC can successfully predict the Follower’s motion from various initial conditions by only relying on online training data, demonstrating its generalizability beyond a priori available training data.
Figure 1: Time-lapse of a lane merge by the GP-MPC. The transparent rectangles show the final predicted positions of vehicles, while for the Follower two rectangles indicate $2\sigma$-bounds on its final predicted position.
@@ VI Conclusions
We presented a Gaussian process-based online learning MPC that uses the dual control effect to learn and predict the dynamics of other agents while accounting for their joint uncertainty. We validated our GP-MPC in a numerical case study on automated lane merging, by leveraging dual MPC to jointly optimize the motion of the ego vehicle and the predictions of a target vehicle through an interactive GP model. By explicitly considering the effect of control inputs on the covariance of predictions of other agents, our GP-MPC can leverage this effect in the tightening of constraints. Through online learning, the GP-MPC can adapt to uncertain and unseen behavior of other agents, showing strong potential for scenarios that extend beyond a priori available training data sets.
The dual control effect is an important notion that can be leveraged by the MPC. While this paper focuses on a proof of concept of the proposed method, the next research steps for learning-based GP-MPC include: (i) incorporating the probing effect of dual control, i.e., active learning, (ii) predicting the residual dynamics of multiple agents, and (iii) providing performance and recursive feasibility guarantees of the MPC. In the context of autonomous driving, planned future work includes validation of the prediction model, the learning of various driving behaviors, and experimental validation.
Figure 2: Average prediction error of Follower’s velocity over the prediction horizon in scenarios in which all planners merge behind (average over 16 experiments, shown in the top figure) and merge in between (average over 20 experiments, shown in the bottom figure).